OpenStudy (anonymous):

Find three consecutive negative integers such that 3 times the product of the first and third exceeds twice the product of the first and second by 32

OpenStudy (anonymous):

Let the first negative integer be = -x second negative integer = -x -1 third negative integer = -x -2 product of first and third = (-x) (-x -2) = x^2 + 2x product of first and second = (-x) (-x -1) = x^2 + x ATQ, 3(x^2 + 2x) - 2(x^2 + x) = 32 3x^2 + 6x -2x^2 - 2x = 32 x^2 + 4x - 32 = 0 x^2 + 8x - 4x - 32 = 0 x(x + 8) -4(x + 8) = 0 (x - 4)(x + 8) = 0 if x - 4 = 0, then x = 4 if x + 8 = 0, then x = -8 so if x = 4 then first negative no is = -x = -4 second negative no is = -x -1 = -5 third negative no is = -x -2 = -6 we cannot use/accept x = -8 because if we take x = -8, then we get positive nos. first no = -x = -(-8) = 8 second no = -x -1 = -(-8) - 1 = 8 - 1 = 7 third no = -x -2 = -(-8) - 2 = 8 -2 = 6 but we have to find negative numbers so -4, -5 and -6 is the right answer